OpenStudy (anonymous):

A 15-g sample of radioactive iodine decays in such a way that the mass remaining after t days is given by m(t) = 15e^-0.087t, where m(t) is measured in grams. After how many days is there inly 5 g remaining? I set 5 = to the equation and it didnt work:(

ganeshie8 (ganeshie8):

thats good start... :) 5 = 15e^(-.087t) 1/3 = e^(-.087t) take ln both sides ln(1/3) = -.087t -1.1 = -.087t t= ?

OpenStudy (anonymous):

So I guess I just forgot to remove 15 by dividing it by 5 again……..stupid I should have looked before I posted this one...

ganeshie8 (ganeshie8):

its okay hehehe :)

OpenStudy (anonymous):

I have another question?!?!

ganeshie8 (ganeshie8):

post

OpenStudy (anonymous):

A man invested $5,000 into an account that pays 8.5% per year compounded quarterly. Find how long it will take for the investment to double…. I just get confused on where to put certain things and exactly what to use.

ganeshie8 (ganeshie8):

Final amount = P ( 1 + r/n)^(n*t)

ganeshie8 (ganeshie8):

you familiar with above formula

ganeshie8 (ganeshie8):

?

OpenStudy (anonymous):

Sorry !! I was working on another problem...

OpenStudy (anonymous):

Yes I have seen that before.

ganeshie8 (ganeshie8):

np \(\huge \textbf{Final amount} = P(1+\frac{r}{n})^{r*t} \) P : initial amount = $5000 r : interest rate per year = .085 n : num. of compounding times per year = 4 t : time in years = ????????????????????????????????????? Final amount : 10,000

ganeshie8 (ganeshie8):

can you find t now

OpenStudy (anonymous):

Okay! thank you so very much!!!

ganeshie8 (ganeshie8):

yw :)

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

Still there I need help with these equations, i must be missing some rule….

ganeshie8 (ganeshie8):

sure ask

OpenStudy (anonymous):

a) 5^x = 4^x+1 b) 2^2x+1 = 3^x-2 c) 50/1+ e^-x = 4 d) log x + log(x-1) = (4x)

ganeshie8 (ganeshie8):

Ok. you wanto understand how to solve x is it

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

23 and 25 says to find the solution...

ganeshie8 (ganeshie8):

hmm who is 23 and 25

OpenStudy (anonymous):

For #23 and #25 Directions say to find the solution..

ganeshie8 (ganeshie8):

those four equation you can solve using only two log properties

OpenStudy (anonymous):

which ones?

ganeshie8 (ganeshie8):

\(\huge 1) \ \log a^m = m\log a\) \(\huge 1) \ \log a + \log b = \log ab\)

OpenStudy (anonymous):

hmm I know those two but I have tried these problems a couple times and I just cant do them!!!

ganeshie8 (ganeshie8):

its okay takes time to understand them. lets try and solve this a) 5^x = 4^x+1

OpenStudy (anonymous):

5 ln x = x ln x+1

OpenStudy (anonymous):

thennn…… divide by 5ln / 4 ln

ganeshie8 (ganeshie8):

wait. let me do the first one showing u step by step. rest three of them u can try ok

OpenStudy (anonymous):

kkkk

ganeshie8 (ganeshie8):

\(5^x = 4^{x+1}\) taking log both sides \(\log(5^x) = \log(4^{x+1})\) \(x \log(5) = (x+1)\log(4)\) \(x \log(5) = x\log(4) + \log(4)\) \(x(\log5 - \log4) = \log(4)\) \(x = \frac{\log 4}{\log 5 - \log 4}\)

ganeshie8 (ganeshie8):

see if it makes sense

OpenStudy (anonymous):

so in step 3, where there is x log (4) + (4) is that from distributing the 4?

ganeshie8 (ganeshie8):

thats from distributing log(4)

OpenStudy (anonymous):

so then in step 4 what happend to the x that was beside the x log (4)

ganeshie8 (ganeshie8):

ok two steps are missing there between step 3 and step 4 3.a ) subtract xlog4 both sides 3.b ) factor the common term x in LHS

OpenStudy (anonymous):

oh okay! that makes sense

ganeshie8 (ganeshie8):

great ! try second one

OpenStudy (anonymous):

kk

ganeshie8 (ganeshie8):

d) is some tricky... but you would be able to figure out... try.... good luck :)

OpenStudy (anonymous):

b) I still cant do :( I guess i get mixed up with steps… the last one i cant do!!!! its hard.

OpenStudy (anonymous):

friend!

ganeshie8 (ganeshie8):

ok that motivates me lol

ganeshie8 (ganeshie8):

\(2^{2x+1} = 3^{x-2}\)

ganeshie8 (ganeshie8):

can i try make u do this ?

OpenStudy (anonymous):

Ive done it!! I would take ln to both side then i get : 3x+1 ln 2 = x-2 ln 3 so then i divided by 2ln both sides and get a different answer!

ganeshie8 (ganeshie8):

thats nice... you understanding logarithms :)

OpenStudy (anonymous):

I get it mostly just get caught up sometimes, gotta give me credit here!

ganeshie8 (ganeshie8):

i see where u getting caught up.. till taking log you are doing it right :) once you bring the exponent to side, you should put () around it and then distribute like this (2x+1) ln 2 = (x-2) ln 3

OpenStudy (anonymous):

ohhhhhhhhhhhhh

ganeshie8 (ganeshie8):

can u distribute nw

OpenStudy (anonymous):

so is it 4x ln = 3xln - 6ln?

OpenStudy (anonymous):

Noppppppee, didnt get the right answer!

ganeshie8 (ganeshie8):

nope let me show u (2x+1) ln 2 = (x-2) ln 3 2x. ln2 + 1. ln2 = x.ln3 -2.ln3

ganeshie8 (ganeshie8):

that "2x" cannot go inside the ln. it has to stay outside lol

ganeshie8 (ganeshie8):

2x. ln2 + 1. ln2 = x.ln3 -2.ln3 subtracting xln3 both side 2x.ln2 - xln3 + 1. ln2 = -2.ln3 subtracting ln2 both sides 2x.ln2 - xln3 = -2.ln3 - ln2 factoring out x x(2ln2 - ln3) = -2ln3 - ln2 x = ?

OpenStudy (anonymous):

whats the 2x. why is there .

ganeshie8 (ganeshie8):

where ? first step ? its there in question itself right

OpenStudy (anonymous):

okay i see nevermind

OpenStudy (anonymous):

next!

ganeshie8 (ganeshie8):

cool.. next is ur turn lol

OpenStudy (anonymous):

I neeed help thats why i am here! trust me ive tried all of these multiple times or else i wouldnt be on here wasting time bugging people to do it for me

ganeshie8 (ganeshie8):

lol its not bugging. you are learning stuff here :)) c) 50/1+ e^-x = 4 \(\huge \frac{50}{1+e^{-x}} = 4\) its like this ?

OpenStudy (anonymous):

yeah

ganeshie8 (ganeshie8):

solutiong is not that straight forward i agree.. .

ganeshie8 (ganeshie8):

multiply both sides with 1+ e^-x

ganeshie8 (ganeshie8):

\(50 = 4(1+e^{-x})\) \(50 = 4+4e^{-x}\) \(46 = 4e^{-x}\) \(46/4 = e^{-x}\) taking ln both sides ln(46/4) = -x x = -ln(46/4)

ganeshie8 (ganeshie8):

it makes sense ?

OpenStudy (anonymous):

yes thank you sorry was doing something else again

ganeshie8 (ganeshie8):

its okay

ganeshie8 (ganeshie8):

last question can u try